Current Position: 15 51.7N 044 16.5W
24 hour progress: 138nm, 5.75kts avg SOG, 1993nm total progress, approximately 970nm to go.
Everyone is doing fine. A bit of stir craziness has set in along with general tiredness. Both boys have been bickering and we tried to get them to nap yesterday which didn't work, unsurprisingly. Cooking dinner was an event as there was a lot of chopping on a rolling boat and it took quite awhile. Then both boys fell asleep before dinner was on the table; however, they woke up on the early side this morning. We will try to make the day interesting as a New Year's Eve treat. There have been daily games of Monopoly and some hangman yesterday. Perhaps today will be a movie marathon and/or more games. Oh, and Jon got to disassemble and rebuild the refrigerator circulator pump again resulting in at least some success.